Find a quadratic equation with roots -1+4i and -1-4i
Answer by ankor@dixie-net.com(22740)

Find a quadratic equation with roots -1+4i and -1-4i
:
x = -1 +/- 4i
x + 1 = +/- 4i
Square both sides
(x+1)^2 = 16(i^2)
FOIL
x^2 + 2x + 1 = -16
x^2 + 2x + 1 + 16 = 0
y = x^2 + 2x + 17 is the equation